the u tube having limbs contains mercury density pm to a level.ifthe left limbs is filled to the top with water pw then the rise of mercury level in the right limb will be

Grade:12th pass

1 Answers

Vikas TU
7 years ago
the pressure at the bottom would be same due to both the limbs.
Hence pressure at bottom => pw*(h-x)*g + pm*g*(x) = p*mg*(h+y)
=> pwgh – pwgx + pmgx = pmgh + pmgy
=> y = (pw*(h-x)*g + pm*g*(x) – pmgh)/(pmg)
is the answer.